### v1

# Baker–Akhiezer specialisation of joint eigenfunctions for hyperbolic relativistic Calogero–Moser Hamiltonians

Abstract
In earlier joint work with Ruijsenaars, we constructed and studied symmetric joint eigenfunctions $J_{N}$ for quantum Hamiltonians of the hyperbolic relativistic $N$-particle Calogero–Moser system. For generic coupling values, they are non-elementary functions that in the $N=2$ case essentially amount to a ‘relativistic’ generalisation of the conical function specialisation of the Gauss hypergeometric function $_{2}F_{1}$. In this paper, we consider a discrete set of coupling values for which the solution to the joint eigenvalue problem is known to be given by functions $\psi_{N}$ of Baker–Akhiezer type, which are elementary, but highly nontrivial, functions. Specifically, we show that $J_{N}$ essentially amounts to the antisymmetrisation of $\psi_{N}$ and, as a byproduct, we obtain a recursive construction of $\psi_{N}$ in terms of an iterated residue formula.
1 Introduction
[…]
$S_{r}(g;x)\equiv\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}f_{-}(g;x_{j}-x_{k})\prod_{l\in I}\exp(-i\hbar\beta\partial_{x_{l}})\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}f_{+}(g;x_{j}-x_{k}),$ (1.1)
where
$f_{\pm}(g;z)=\left(\frac{\sinh(\mu(z\pm i\beta g)/2)}{\sinh(\mu z/2)}\right)^{1/2},$ (1.2)
is of particular importance. Here, $r=1,\ldots,N$, it is natural to view $\beta>0$ as $1/mc$, with $m$ the particle mass and $c$ the speed of light, and in the nonrelativistic limit $c\to\infty$ pairwise commuting Hamiltonians of the nonrelativistic hyperbolic Calogero–Moser system are recovered; see e.g. the surveys [Rui99, Hal25] and references therein.
[…]
Further fundamental properties of the joint eigenfunctions were obtained by Belousov et al. in a number of recent papers [BDKK24a, BDKK24b, BDKK24c, BDKK24d]. Their results include integral equations, a reflection symmetry of the coupling constant, self-duality under interchange of geometric and spectral variables as well as orthogonality and completeness relations.
[…]
$D^{r}_{N}(z;q,t)\equiv t^{r(r-1)/2}\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{tx_{j}-x_{k}}{x_{j}-x_{k}}\prod_{l\in I}T_{q,x_{l}},$ (1.3)
where $(T_{q,x_{l}}f)(x_{1},\ldots,x_{k},\ldots,x_{N})=f(x_{1},\ldots,qx_{l},\ldots,x_{N})$. (The precise relationship can be gleaned from (2.6) and (3.4).) They act on the space of symmetric polynomials in $N$ variables, on which they are simultaneously diagonalised by the symmetric $GL_{N}$ type Macdonald polynomials; see e.g. [Mac95, Sto21]. For parameter values of the form $t=q^{m}$, $m\in{\mathbb{Z}}$, Etingof and Styrkas [ES98] and Chalykh [Cha02] constructed and studied non-symmetric joint eigenfunctions of Baker–Akhiezer (BA) type and obtained, in particular, a generalized Weyl character formula for the Macdonald polynomials, first conjectured by Felder and Varchenko [FV97], where the BA-function replaces the exponential function.
[…]
For the coupling values $g=m\hbar$, $m\in\mathbb{Z}_{+}$, under consideration, our main result provides an expansion of symmetric joint eigenfunctions in terms of eigenfunctions that are ‘asymptotically free’ deep inside Weyl chambers $x_{\sigma(1)}>x_{\sigma(2)}>\cdots>x_{\sigma(N)}$, $\sigma\in S_{N}$. In the case of the Macdonald operators and generic parameter values, such asymptotically free solutions, often referred to as (q-)Harisch–Chandra series, have been studied in detail by Letzter and Stokman [LS08], van Meer and Stokman [vMS10] as well as Noumi and Shiraishi [NS12]; and Stokman [Sto14] derived a corresponding ($c$-function) expansion of Cherednik’s basic hypergeometric function associated to root systems. To obtain an analogous expansion for the joint $S_{r}$-eigenfunction from [HR14] in the case of generic $g$-values is, I believe, an interesting open problem to which I hope to return elsewhere. In the $N=2$ case, the first steps in this direction can be found in a recent paper by di Francesco et al. [FKKSS24], in which they obtained a number of interesting results on these eigenfunction, including how Harish–Chandra series are recovered by suitably closing contours in one of their integral representations and computing residues.
The remainder of the paper is structured as follows: In Section 2, we briefly recall definitions and properties pertaining to joint eigenfunctions of the Ruijsenaars operators (1.1) and BA-(eigen)functions for the Macdonald operators (1.3); and, in Section 3, we give the precise formulations and proofs of our results.
2 Preliminaries
2.1 Joint eigenfunctions
[…]
$a_{+}\equiv 2\pi/\mu>0,\quad\mathrm{(imaginary~period/interaction~length)}$ (2.1)
and
$a_{-}\equiv\hbar\beta>0,\quad\mathrm{(shift~step~size/Compton~wavelength)}$ (2.2)
and replace the coupling $g$ by the parameter
$b\equiv\beta g.$ (2.3)
[…]
$W_{N}(b;x)\equiv\prod_{1\leq j<k\leq N}\prod_{\delta=+,-}\frac{G(\delta(x_{j}-x_{k})+i(a_{+}+a_{-})/2)}{G(\delta(x_{j}-x_{k})+i(a_{+}+a_{-})/2-ib)},$ (2.4)
where $G(z)\equiv G(a_{+},a_{-};z)$ is the hyperbolic Gamma function from [Rui97], which is $a_{+}\leftrightarrow a_{-}$ invariant, meromorphic in $z$ and satisfies the analytic difference equations
$\frac{G(z+ia_{\delta}/2)}{G(z-ia_{\delta}/2)}=2\cosh(\pi z/a_{-\delta}),\ \ \ \delta=+,-;$ (2.5)
see loc. cit. Prop. III.1 and Prop. III.2. Indeed, using these difference equations, it is readily verified that
$\begin{split}A_{r,\delta}(x)&\equiv W_{N}(x)^{-1/2}H_{r,\delta}(b;x)W_{N}(x)^{1/2}\\ &=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh(\pi(x_{j}-x_{k}-ib)/a_{\delta})}{\sinh(\pi(x_{j}-x_{k})/a_{\delta})}\prod_{l\in I}\exp(-ia_{-\delta}\partial_{x_{l}}),\end{split}$ (2.6)
which, in contrast to $H_{r,\delta}$, act on the space of meromorphic functions. (We recall that, for $(a_{+},a_{-},b)\in(0,\infty)^{3}$ such that $b<a_{+}+a_{-}$ and $x\in{\mathbb{R}}^{N}$, the weight function is regular and positive.)
In [HR14] functions $J_{N}(b;x,y)$ having the joint eigenfunction property
$A_{r,\delta}J_{N}(x,y)=S_{r}(e^{\pi y_{1}/a_{\delta}},\ldots,e^{\pi y_{N}/a_{\delta}})J_{N}(x,y),$ (2.7)
with the elementary symmetric functions
$S_{r}(t)\equiv\sum_{1\leq j_{1}<\cdots j_{r}\leq N}t_{j_{1}}\cdots t_{j_{r}},$ (2.8)
are constructed recursively by an explicit integral formula, with integrand built from $J_{N-1}$, the weight function $W_{N-1}$ as well as the kernel function
${\cal S}_{N}^{\sharp}(b;x,y)\equiv\prod_{j=1}^{N}\prod_{k=1}^{N-1}\frac{G(x_{j}-y_{k}-ib/2)}{G(x_{j}-y_{k}+ib/2)}.$ (2.9)
More precisely, with $J_{1}(x,y)\equiv\exp(2\pi ixy/a_{+}a_{-})$ as the starting point for the recursion, $J_{N}$, $N>1$, is given by
$J_{N}(b;x,y)=\frac{\exp\left(\frac{2\pi i}{a_{+}a_{-}}y_{N}\sum_{j=1}^{N}x_{j}\right)}{(N-1)!}\\ \cdot\int_{{\mathbb{R}}^{N-1}}dzW_{N-1}(b;z){\cal S}^{\sharp}_{N}(b;x,y)J_{N-1}(b;z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$ (2.10)
From Prop. III.2 in [Rui97], we recall that the hyperbolic gamma function is scale invariant, in the sense that
$G(\lambda a_{+},\lambda a_{-};\lambda z)=G(a_{+},a_{-};z),\quad\lambda\in(0,\infty).$ (2.11)
Since $J_{N}$ is constructed almost entirely from $G(z)$, it is easily seen to have the invariance property
$J_{N}(\lambda a_{+},\lambda a_{-},\lambda b;\lambda x,\lambda y)=\lambda^{N(N-1)/2}J_{N}(a_{+},a_{-},b;x,y).$ (2.12)
Hence, we may and shall restrict attention to
$\min(a_{+},a_{-})=1,\ \ \ \max(a_{+},a_{-})\equiv a\geq 1,$ (2.13)
without any loss of generality. Moreover, to facilitate the comparison with the pertinent BA-function, it is expedient to renormalize $J_{N}$ and introduce the function
$\Phi_{N}(a,b;x,y)\equiv\left(\frac{G(1,a;ib-i(1+a)/2)}{\sqrt{a}}\right)^{N-1}J_{N}(1,a,b;x,y),$ (2.14)
which, in particular, satisfies the simple duality relation
$\Phi_{N}(b;x,y)=\Phi_{N}(1+a-b;y,x),$ (2.15)
as conjectured in [HR14], verified for $N\leq 3$ in [HR18a] and proven for arbitrary $N$ in [BDKK24b] (see Thm. 5 and Eqs. (1.53)–(1.54)).
2.2 Baker–Akhiezer functions
Take $q\in\mathbb{C}^{\times}\equiv{\mathbb{C}}\setminus\{0\}$, $m\in\mathbb{Z}_{+}$ and let $\psi_{N}$ be a function of $x,y\in\mathbb{C}^{N}$ that is of the form
$\psi_{N}(x,y)=q^{2(x,y)}\sum_{\nu}\psi_{N,\nu}(x)q^{2(\nu,y)},$ (2.16)
where the sum extends over weight vectors
$\nu=\sum_{1\leq j<k\leq N}\left(\frac{m}{2}-l_{jk}\right)(e_{j}-e_{k}),\ \ \ l_{jk}=0,\ldots,m.$ (2.17)
Suppose, in addition, that $\psi_{N}$ satisfies the vanishing condition
$\psi_{N}(x,y+s(e_{j}-e_{k})/2)-\psi_{N}(x,y-s(e_{j}-e_{k})/2)=0,\ \ \ q^{2(y_{j}-y_{k})}=1,$ (2.18)
for all $1\leq j<k\leq N$ and $s=1,\ldots,m$. Then it is called a Baker–Akhiezer (BA) function associated with the root system $A_{N-1}$ and (positive) integer parameter $m$. Existence is established by different constructions in Sect. 5 of [ES98] and Sect. 3.2 of [Cha02]. We note that the iterated residue formula obtained in this paper yields an additional (constructive) existence proof.
Assuming $q$ is not a root of unity, we know from [ES98, Cha02] that a function $\psi_{N}$ with the above properties is unique up to a choice of normalisation; and that, imposing the normalisation condition
$\psi_{N,\rho_{N}}(x)=\prod_{1\leq j<k\leq N}\prod_{j=1}^{m}[j+x_{k}-x_{j}],\ \ \ [z]\equiv q^{z}-q^{-z},$ (2.19)
yields a self-dual BA function, in the sense that $\psi_{N}(x,y)=\psi_{N}(y,x)$. Here, we have used the standard notation $\rho_{N}\equiv\rho_{N}(m)$ with
$\rho_{N}(m)\equiv\frac{m}{2}\sum_{1\leq j<k\leq N}(e_{j}-e_{k})=\frac{m}{2}\sum_{j=1}^{N}(N-2j+1)e_{j},$ (2.20)
obtained by setting all $l_{ij}=0$ in (2.17). The self-dual BA-function satisfies the bispectral system of $q$-difference equations
$D^{r}_{N}(q^{2y};q^{2},q^{-2m})\psi(x,y)$ $=S_{r}(q^{2x})\psi(x,y),$ (2.21)
$D^{r}_{N}(q^{2x};q^{2},q^{-2m})\psi(x,y)$ $=S_{r}(q^{2y})\psi(x,y),$ (2.22)
with $r=1,\ldots,N$, $q^{2x}\equiv(q^{2x_{1}},\ldots,q^{2x_{N}})$ and $q^{2y}\equiv(q^{2y_{1}},\ldots,q^{2y_{N}})$.
3 Baker–Akhiezer specialisation
Taking $m\in\mathbb{Z}_{+}$ and letting
$\mathcal{\delta}_{N}(x;m)=\prod_{1\leq j<k\leq N}\prod_{n=-m}^{m}[x_{j}-x_{k}+n],$ (3.1)
we recall the well-known similarity transform
$\mathcal{\delta}_{N}(x;m-1)^{-1}D^{r}_{N}(q^{2x};q^{2},q^{-2(m-1)})\mathcal{\delta}_{N}(x;m-1)=D^{r}_{N}(q^{2x};q^{2},q^{2m}),$ (3.2)
which is easily verified by a direct computation; and if we restrict attention to $|q|=1$ using the parametrisation
$q=e^{-i\pi/a},\ \ \ a\in{\mathbb{R}},$ (3.3)
then it is readily seen that
$D^{r}_{N}(q^{2ix};q^{2},q^{2m})=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}-im)\big)}{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k})\big)}\prod_{l\in I}\exp(-i\partial_{x_{l}}).$ (3.4)
[…]
Theorem 3.1.
Let $m\in\mathbb{Z}_{+}$ and let a>m1a>m-1 be such that $\exp(-i\pi/a)$ is not a root of unity. Then, we have
$\Phi_{N}(a,m;x,y)=C_{N}\frac{\sum_{\sigma\in S_{N}}(-)^{\sigma}\psi_{N}(\exp(-i\pi/a),m-1;i\sigma(x),iy)}{\prod_{1\leq j<k\leq N}2\sinh\big(\frac{\pi}{a}(y_{k}-y_{j})\big)\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}$ (3.5)
with the constant $C_{N}\equiv C_{N}(a,m)$ given by
$C_{N}=(2i^{m})^{N(N-1)/2}\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{(N-1)(N-2)/2}.$ (3.6)
Before proceeding to our proof of the theorem, we note that the weight function $W_{N}$ (2.4) and kernel function ${\cal S}^{\sharp}_{N}$ (2.9) reduce to elementary functions when $b=m\in\mathbb{Z}_{+}$. More precisely, from the difference equation (2.5) satisfied by $G(z)$, we see that
$W_{N}(1,a,m;x)=\prod_{1\leq j\neq k\leq N}\prod_{n=0}^{m-1}2\sinh\left(\frac{\pi}{a}(x_{j}-x_{k}-in)\right)$ (3.7)
and
${\cal S}^{\sharp}_{N}(1,a,m;x,y)=\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}\frac{1}{2\cosh\big(\frac{\pi}{a}(x_{j}-y_{k}+i(m-2n-1)/2)\big)}.$ (3.8)
We note that, up to an overall numerical factor, $W_{N}(1,a,m;x)$ equals the specialisation $\Delta(q^{2ix};q^{2},q^{2m})$ of Macdonald’s weight function $\Delta$, while ${\cal S}^{\sharp}_{N}(1,a,m;x,y)$ is obtained by a similar specialisation of his product function $\Pi$ after removing a factor depending only on $\sum_{j=1}^{N}(x_{j}-y_{j})$; cf. Eqs. (9.2) and (2.5), respectively, in Section VI of [Mac95]. Furthermore, (2.5) and the formula $G(1,a;i(1-a)/2)=1/\sqrt{a}$ (see Eq. (3.38) in [Rui97]) entail that
$\frac{G(1,a;im-i(1+a)/2)}{\sqrt{a}}=\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}.$ (3.9)
Consequently, the $b=m$ specialisation of $\Phi_{N}$ is given by recursively by
$\Phi_{N}(x,y)=\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a(N-1)!}e^{\frac{2\pi i}{a}y_{N}(x_{1}+\cdots+x_{N})}\\ \cdot\int_{\mathbb{R}^{N-1}}dz\frac{\Phi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j\neq k\leq N-1}\prod_{n=0}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{j}-z_{k}-in)\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\cosh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)}$ (3.10)
with $\Phi_{1}(x,y)=\exp\left(\frac{2\pi}{a}xy\right)$.
As it turns out, the above integrand is a meromorphic function whose poles are due solely to the $\cosh$-factors in the denominator. The first main observation in our proof of Thm. 3.1 is that $\Phi_{N}(x,y)\prod_{1\leq j<k\leq N}2\sinh(\pi(y_{j}-y_{k}))$ can be rewritten as the symmetrisation of a function $\varphi_{N}(x,y)$, given recursively by an iterated residue formula.
More specifically, the formula involves denominator factors $\sinh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)$, producing poles at
$z_{k}=x_{j}+i(m-2n-1)/2+i\ell a,\ \ \ n=0,\ldots,m-1,\ \ \ell\in\mathbb{Z},$ (3.11)
where $j=1,\ldots,N$ and $k=1,\ldots,N-1$. Assuming that ${\rm Re}\,(x_{j}-x_{k})\neq 0$ for all $1\leq j<k\leq N$ and a>m1a>m-1, we can find contours $\gamma_{k}$ that only encircle the poles corresponding to a fixed value of $k=1,\ldots,N-1$ and $\ell=0$; see the discussions preceding Props. 3.3 and 3.6 for explicit examples. Up to normalisation, the recursive formula defining $\varphi_{N}$ can then be obtained from (3.10) by the substitutions $\Phi\to\varphi$, ${\mathbb{R}}^{N-1}\to\gamma_{1}\times\cdots\times\gamma_{N-1}$ and $\cosh\to\sinh$.
The second main observation, which more or less completes the proof of Thm. 3.1, is that $\varphi_{N}$ essentially amounts to a renormalisation of $\psi_{N}$ and, as a byproduct, we thus find the following recursive construction of the self-dual BA-function $\psi_{N}$.
Proposition 3.2.
Let $q=e^{-i\pi/a}$. Then, under the above assumptions on $m$ and $a$, the BA-function $\psi_{N}$ is obtained from $\psi_{N-1}$ by the iterated residue formula
$\psi_{N}(ix,iy)\\ =\left(\frac{\prod_{n=1}^{m}\sin(\pi n/a)}{2ai^{m+1}}\right)^{N-1}\frac{W_{N}(1,a,m+1;z)}{\prod_{1\leq j<k\leq N}2\sinh\left(\frac{\pi}{a}(x_{k}-x_{j})\right)}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\frac{\psi_{N-1}(iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j<k\leq N-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n)\big)\big)},$ (3.12)
where $\underline{\gamma}\equiv\gamma_{1}\times\cdots\times\gamma_{N-1}$.
This result provides a natural generalisation of the iterated residue formulae for BA-functions for rational and trigonometric Calogero–Moser–Sutherland operators obtained by Felder and Veselov [FV09].
In the two subsections below, we provide our proofs of Thm. 3.1 and Prop. 3.2. To begin with, we establish the $N=2$ cases of the results. These we then use as the basis for proofs by induction on $N$.
3.1 The $N=2$ case
Setting $N=2$ in (3.10), we deduce
$\frac{2a\sinh\big(\frac{\pi}{a}(y_{1}-y_{2})\big)}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\Phi_{2}(x,y)\\ =\int_{\mathbb{R}}dz\frac{e^{\frac{2\pi i}{a}(y_{1}-y_{2})(z-ia/2)}-e^{\frac{2\pi i}{a}(y_{1}-y_{2})(z+ia/2)}}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\cosh\big(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2)\big)}\\ =(-1)^{m}\left(\int_{{\mathbb{R}}-ia/2}dz-\int_{{\mathbb{R}}+ia/2}dz\right)\frac{e^{\frac{2\pi i}{a}(y_{1}-y_{2})z}}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2)\big)}.$ (3.13)
We note that a>m1a>m-1 and $x,y\in\mathbb{R}^{2}$ ensures that $\Phi_{2}$ is well-defined and that the only poles of the integrand located within the strip $|{\rm Im}\,z|<a/2$ are
$z=x_{j}+i(m-2n-1)/2,\ \ \ j=1,2,\ \ n=0,\ldots,m-1.$ (3.14)
At first, we assume $x=(x_{1},x_{2})\in{\mathbb{R}}^{2}$ is such that $x_{1}\neq x_{2}$, which ensures that we can introduce a contour $\gamma_{1}$ encircling only the poles (3.14) with $j=1$ counterclockwise. For example, with $\delta\equiv|x_{1}-x_{2}|>0$, we can take the contour consisting of line segments $x_{1}+u\pm ia/2$, $|u|\leq\delta/2$, and $x_{1}\pm\delta/2+iv$, $|v|\leq a/2$. Introducing the function
$\varphi_{2}(x,y)\equiv e^{\frac{2\pi i}{a}y_{2}(x_{1}+x_{2})}\int_{\gamma_{1}}dz\frac{e^{\frac{2\pi i}{a}(y_{1}-y_{2})z}}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2)\big)},$ (3.15)
the following result is an easy consequence of (3.13) and Cauchy’s residue theorem.
Proposition 3.3.
For $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
$\Phi_{2}(x,y)=(-1)^{m}\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\frac{\sum_{\sigma\in S_{2}}\varphi_{2}(\sigma(x),y)}{2\sinh\big(\frac{\pi}{a}(y_{1}-y_{2})\big)}.$ (3.16)
We proceed to establish the precise connection between $\varphi_{2}$ and the self-dual BA-function $\psi_{2}$. As detailed in the following proposition, when applying Cauchy’s residue thm. to (3.15) we obtain a series expansion for $\varphi_{2}$ analogous to the expansion of $\psi_{2}$ given by the $N=2$ instance of (2.16).
Proposition 3.4.
Assuming $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
$\varphi_{2}(x,y)=e^{\frac{2\pi i}{a}(x,y)}\sum_{l=0}^{m-1}\varphi_{2,l}(x)e^{\frac{2\pi}{a}\left(\frac{m-1}{2}-l\right)(y_{1}-y_{2})}$ (3.17)
with
$\varphi_{2,l}(x)=\frac{2a(-i)^{m}}{\prod_{n\neq l}2\sin\big(\frac{\pi}{a}(n-l)\big)}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))\big)}.$ (3.18)
Next, we show that $\varphi_{2}$ has vanishing properties similar to those of $\psi_{2}$ given by (2.18) with $N=2$.
Proposition 3.5.
For $s=1,\ldots,m-1$, we have
$\varphi_{2}\left(x,y+is(e_{1}-e_{2})/2\right)-\varphi_{2}\left(x,y-is(e_{1}-e_{2})/2\right)=0,\ \ \ e^{\frac{2\pi}{a}(y_{1}-y_{2})}=1.$ (3.19)
Proof.
Up to an overall factor $2\exp(\frac{2\pi i}{a}(y_{2}(x_{1}+x_{2})))$, the LHS of (3.19) is given by the integral
$\int_{\gamma_{1}}dz\frac{\exp\big(\frac{2\pi i}{a}(y_{1}-y_{2})z\big)\sinh(\frac{\pi s}{a}(x_{1}+x_{2}-2z))}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2))},$ (3.20)
which, by Cauchy’s theorem, equals
$(-i)^{m}e^{\frac{2\pi i}{a}(y_{1}-y_{2})x_{1}}\sum_{l=0}^{m-1}\frac{2a}{\prod_{n\neq l}\sin(\frac{\pi}{a}(n-l))}\\ \cdot\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}-i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}e^{-\frac{\pi}{a}(m-2l-1)(y_{1}-y_{2})}.$ (3.21)
Since $e^{-\frac{\pi}{a}(m-2l-1)(y_{1}-y_{2})}=e^{\frac{\pi}{a}(m-1)(y_{1}-y_{2})}$ when $e^{\frac{2\pi}{a}(y_{1}-y_{2})}=1$, we only have to show that
$\sum_{l=0}^{m-1}\frac{2a}{\prod_{n\neq l}\sin(\frac{\pi}{a}(n-l))}\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}-i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}=0.$ (3.22)
To this end, we note that $s<m$ entails that the left-hand side is a meromorphic function in $x_{1}-x_{2}$ that is bounded away from its poles and decays exponentially as $|{\rm Re}\,(x_{1}-x_{2})|\to\infty$. Hence, Liouville’s theorem will imply the above identity once we show that the residues at the (simple) poles
$x_{2}-x_{1}+i\ell=i\ell^{\prime}a,\ \ \ \ell=-m+1,\ldots,m-1,\ \ \ell^{\prime}\in\mathbb{Z},$ (3.23)
all vanish; and, by iaia-(anti)periodicity, we may and shall restrict attention to $\ell^{\prime}=0$. For a fixed $\ell=-m+1,\ldots,m-1$, we observe that there exists $n=0,\ldots,m-1$ such that $n-l=\ell$ if and only if
$l=\max(-\ell,0),\ldots,\min(m-1,m-1-\ell).$ (3.24)
It follows that the residue of the left-hand side of (3.22) at $x_{2}-x_{1}+i\ell=0$ is proportional to
$\sum_{l=\max(-\ell,0)}^{\min(m-1,m-1-\ell)}f_{l},\ \ \ f_{l}(a)\equiv\frac{2a}{\prod_{n\neq l}\sin(\frac{\pi}{a}(n-l))}\frac{\sin(\frac{\pi s}{a}(2l+\ell-m+1))}{\prod_{n\neq l+\ell}\sin(\frac{\pi}{a}(n-l-\ell))}.$ (3.25)
On the set of integers (3.24), we have the involution $\sigma:l\mapsto m-1-\ell-l$. Since $f_{\sigma(l)}=-f_{l}$, as is easily checked, this clearly implies that the residue vanishes. ∎
Setting $l=0$ in (3.18), we find that
$\varphi_{2,0}(x)=\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}.$ (3.26)
Comparing (3.26), (3.19) and (3.17) with the $N=2$ instances of (2.19), (2.18) and (2.16), respectively, the uniqueness of the BA-function, once a normalisation has been fixed, implies that
$\varphi_{2}(x,y)=\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\frac{\psi_{N}(\exp(-i\pi/a),m-1;i\sigma(x),iy)}{\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}.$ (3.27)
When combined with Prop. 3.3, we obtain the $N=2$ case of Thm. 3.1; and a comparison with (3.15) yields the $N=2$ instance of Prop. 3.2.
3.2 The inductive step
We can now use the $N=2$ case as the basis for proofs by induction. More precisely, we consider $N\geq 3$ and assume that Thm. 3.1 along with Props. 3.63.10 hold true when $N$ is replaced by N1N-1. (In the $N=2$ case, the pertinent results have all been established above.)
To begin with, we require that $x=(x_{1},\ldots,x_{N})\in{\mathbb{R}}^{N}$ with $x_{j}\neq x_{k}$ for all $1\leq j<k\leq N$. Letting $\delta=\min\{|x_{j}-x_{k}|\mid 1\leq j<k\leq N\}$, we construct $N$ contours $\gamma_{j}$, $j=1,\ldots,N$, from the line segments $x_{j}+u\pm ia/2$, $|u|\leq\delta/2$, and $x_{j}\pm\delta/2+iv$, $|v|\leq a/2$; and define the function
$\varphi_{N}(x,y)\equiv e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\frac{\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j\neq k\leq N-1}\prod_{n=0}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{j}-z_{k}-in)\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.28)
with $\underline{\gamma}\equiv\gamma_{1}\times\cdots\times\gamma_{N-1}$, and where the $z_{j}$-contour $\gamma_{j}$ only encircles the simple poles
$z_{j}=x_{j}+i(m-2n-1)/2,\ \ \ n=0,\ldots,m-1.$ (3.29)
The arbitrary-$N$ generalisation of Prop. 3.3 now follows.
Proposition 3.6.
Assuming $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we get
$\prod_{1\leq j<k\leq N}2\sinh(\pi(y_{j}-y_{k}))\cdot\Phi_{N}(x,y)\\ =(-1)^{mN(N-1)/2}\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}\sum_{\sigma\in S_{N}}\varphi_{N}(\sigma(x),y).$ (3.30)
Proof.
To ease the notation, we suppress dependence on the parameters $a$ and $m$ throughout the proof. Substituting (3.10) in the left-hand side of (3.30), using (3.7)–(3.8) to reduce the size of the resulting expression and invoking the above statement after taking $N\to N-1$, we obtain
$\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}\prod_{j=1}^{N-1}2\sinh(\pi(y_{j}-y_{N}))\frac{e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}}{(N-1)!}\\ \cdot\sum_{\sigma\in S_{N-1}}\int_{\mathbb{R}^{N-1}}dzW_{N-1}(z){\cal S}^{\sharp}_{N}(x,z)\varphi_{N-1}(\sigma(z),(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$
Thanks to the $S_{N-1}$-invariance of the domain of integration $\mathbb{R}^{N-1}$ as well as $W_{N-1}(z)$ and ${\cal S}^{\sharp}_{N}(x,z)$, we can replace the sum over $\sigma\in S_{N-1}$ by a factor (N1)!(N-1)!. By also substituting $2\sinh(\pi(y_{j}-y_{N}))=e^{\pi(y_{j}-y_{N})}-e^{-\pi(y_{j}-y_{N})}$, we thus obtain
$\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\sum_{\delta\in\{\pm 1\}^{N-1}}\delta_{1}\cdots\delta_{N-1}\\ \cdot\int_{\mathbb{R}^{N-1}}dzW_{N-1}(z){\cal S}^{\sharp}_{N}(x,z)e^{\pi\sum_{j=1}^{N-1}\delta_{j}(y_{j}-y_{N})}\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$ (3.31)
To rewrite the integral further we shall make use of the invariance property
$e^{\pi\sum_{j=1}^{N-1}\delta_{j}u_{j}}\varphi_{N-1}(z,u)=\varphi_{N-1}(z-(ia/2)\delta,u)\prod_{j=1}^{N-1}\delta_{j}^{m(N-2)}.$ (3.32)
When N1=2N-1=2 it is a simple consequence of Prop. 3.4 and, using this as the basis for an induction argument, the validity of the formula in the general-$N$ case is readily inferred from (3.41)–(3.42) and the observation that $W_{N-1}(z+(ia/2)\delta)=W_{N-1}(z)$. Substituting (3.32) (with $u_{j}=y_{j}-y_{N}$) in the above integral, taking $z\to z+(ia/2)\delta$, under which $W_{N-1}(z)$ is invariant, and noting that
${\cal S}^{\sharp}_{N}(x,z+(ia/2)\delta)=\frac{(-1)^{mN(N-1)/2}\prod_{j=1}^{N-1}\delta_{j}^{mN}}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)},$ (3.33)
we arrive at
$(-1)^{mN(N-1)/2}\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\sum_{\delta\in\{\pm 1\}^{N-1}}\delta_{1}\cdots\delta_{N-1}\\ \cdot\int_{{\mathbb{R}}-\frac{ia}{2}\delta_{1}}dz_{1}\cdots\int_{{\mathbb{R}}-\frac{ia}{2}\delta_{N-1}}dz_{N-1}\frac{W_{N-1}(z)\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.34)
where the sum of integrals amounts to
$\int_{{\cal C}^{N-1}}dz\frac{W_{N-1}(z)\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)}$ (3.35)
with $\cal C$} the contour consisting of the lines ${\mathbb{R}}\pm ia/2$, traversed from right/left to left/right. By Cauchy’s residue theorem, this one integral equals
$\sum_{\sigma\in S_{N}}\int_{\gamma_{\sigma(1)}}dz_{1}\cdots\int_{\gamma_{\sigma(N-1)}}dz_{N-1}\\ \cdot\frac{W_{N-1}(z)\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)}+R(x,y)$
for some remainder term $R(x,y)$, resulting from residues of the integrand at points $z=(z_{1},\ldots,z_{N-1})$ given by
$z_{j}=x_{i_{j}}+i(m-2n_{j}+1)/2,\ \ \ j=1,\ldots,N-1,$ (3.36)
where at least two $i_{j}$ coincide.
Referring back to (3.28), we find that it remains to prove that $R=0$. To this end, we take $N\to N-1$ in Prop. 3.10 and use the resulting formula to rewrite the integrand as
$C_{N-1}\frac{\psi_{N-1}(e^{-i\pi/a},m-1;iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j<k\leq N-1}\sinh\big(\frac{\pi}{a}(z_{j}-z_{k})\big)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.37)
where $C_{N-1}$ is a constant (for fixed $m$ and $a$), whose value is of no consequence for the arguments below. From Prop. 4.4 in [Cha02] (part (iii)), we know that $\psi_{N-1}(u,v)$ is an entire function. The presence of the factors $\sinh\big(\frac{\pi}{a}(z_{j}-z_{k})\big)$ thus entails that (iterated) residues at points (3.36) such that $i_{j}=i_{k}$ and $n_{j}=n_{k}$ for some $1\leq j\neq k\leq N-1$ all vanish. In the remaining cases we have $n_{j}\neq n_{k}$ while $i_{j}=i_{k}$. Thanks to the $S_{N-1}$-invariance of the integrand (cf. Lemma 3.8), we may and shall restrict attention to $j=1$ and $k=2$. Letting $\widetilde{\varphi}_{N-1}(z)=\psi_{N-1}(e^{-i\pi/a},m-1;iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))$, $t=x_{i_{1}}=x_{i_{2}}$ and $s=n_{1}-n_{2}$, we observe that
$\tilde{\varphi}_{N-1}\big((t+i(m-2n_{1}-1)/2,t+i(m-2n_{2}-1)/2,\ldots)\big)\\ =\tilde{\varphi}_{N-1}\big((t+i(m-n_{1}-n_{2}-1)/2-is/2,t+i(m-n_{1}-n_{2}-1)/2+is/2,\ldots)\big).$
Note that the self-duality of $\psi_{N-1}$ entails that $\widetilde{\varphi}_{N-1}(z)$ satisfies the vanishing conditions in Prop. 3.9 with $N\to N-1$ in $z$. Since each $n_{j}=0,\ldots,m-1$, we have $s=-m+1,\ldots m-1$. From the vanishing conditions for $\tilde{\varphi}$, we can thus infer that its specialisation at (3.36) with $i_{1}=i_{2}$ is invariant under the interchange $n_{1}\leftrightarrow n_{2}$. Moreover, we have
$\sinh(\pi(z_{1}-z_{2})/a)=\sinh(i\pi(n_{2}-n_{1})/a),$ (3.38)
so that the same specialisation of $\varphi(z)W_{N-1}(z)$ is antisymmetric under $n_{1}\leftrightarrow n_{2}$. In fact, more generally, when $i_{1}=\cdots=i_{M}$, $M\geq 2$, coincide, we can, similarly, show that antisymmetry extends to any permutation of the corresponding $M$ distinct integers $n_{1},\ldots,n_{M}$. (Essentially, we need only to decompose such a permutation in terms of (elementary) transpositions and appeal to the above discussion.) Either way, we see that the residues corresponding to a fixed choice of indices $i_{j}$, at least two of which coincide, and integers $n_{j}$, fixed up to permutations among the ones corresponding to equal $i_{j}$, cancel (in pairs) and consequently that $R=0$. ∎
We proceed to show that $\varphi_{N}$ (3.28) essentially amounts to the BA-function reviewed above. This requires that we establish two things: the relevant series expansion and vanishing properties.
Proposition 3.7.
We have
$\varphi_{N}(x;y)=e^{\frac{2\pi i}{a}(x,y)}\sum_{\nu}\varphi_{N,\nu}(x)e^{\frac{2\pi}{a}(\nu,y)}$ (3.39)
with the sum running over weight vectors of the form (2.17) with $m\to m-1$ and where
$\begin{split}\varphi_{N,\rho_{N}(m-1)}(x)&=\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{1}{\prod_{1\leq j<k\leq N}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}.\end{split}$ (3.40)
Proof.
By our induction assumption, we may invoke the proposition after taking $N\to N-1$. Substituting the resulting expansion (3.39) in (3.28) and making use of Cauchy’s residue theorem, we deduce
$\varphi_{N}(x;y)=e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\sum_{\nu^{\prime}}e^{\frac{2\pi}{a}(\nu^{\prime},y)}\\ \cdot\int_{\underline{\gamma}}dz\frac{W_{N-1}(z)\varphi_{N-1,\nu^{\prime}}(z)e^{\frac{2\pi i}{a}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)}\\ =e^{\frac{2\pi i}{a}(x,y)}\sum_{\nu^{\prime}}e^{\frac{2\pi}{a}(\nu^{\prime},(y_{1},\ldots,y_{N-1}))}\sum_{n_{1}=0}^{m-1}\cdots\sum_{n_{N-1}=0}^{m-1}C_{\nu^{\prime};n_{1},\ldots,n_{N-1}}(x)e^{\frac{2\pi}{a}\sum_{k=1}^{N-1}(n_{k}-(m-1)/2)(y_{k}-y_{N})},$ (3.41)
with $\nu^{\prime}$ running over the set of weight vectors obtained after taking $N\to N-1$ and $m\to m-1$ in (2.17), and where
$C_{v^{\prime};n_{1},\ldots,n_{N-1}}(x)\\ =(-2ia)^{N-1}\frac{W_{N-1}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{n^{\prime}\neq n_{k}}2\sinh\big(\frac{i\pi}{a}(n_{k}-n^{\prime})\big)}\\ \cdot\frac{\varphi_{N-1,\nu^{\prime}}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(n_{k}-n^{\prime})\big)}$ (3.42)
and we have used that $(\nu^{\prime},(y_{N},\ldots,y_{N}))=0$, since $\nu^{\prime}$ is a linear combination of vectors $e_{j}-e_{k}$ with $1\leq j<k\leq N-1$. Setting $l_{j,N-1}=m-1-n_{j}=0,\ldots,m-1$, we get the equality
$\sum_{1\leq j<k\leq N-1}\left(\frac{m-1}{2}-l_{jk}\right)(e_{j}-e_{k})+\sum_{j=1}^{N-1}\left(n_{j}-\frac{m-1}{2}\right)(e_{j}-e_{N})\\ =\sum_{1\leq j<k\leq N}\left(\frac{m-1}{2}-l_{jk}\right)(e_{j}-e_{k})$ (3.43)
and the validity of (3.39) with $\nu$ from (2.17) clearly follows.
Since the only term in (3.41) that contributes to $\varphi_{N,\rho_{N}(m-1)}(x)$ corresponds to $\nu^{\prime}=\rho_{N-1}(m-1)$ and $n_{1}=\cdots=n_{N-1}=m-1$, we have
$\varphi_{N,\rho_{N}(m-1)}(x)=\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a\big)}\right)^{N-1}\\ \cdot\frac{W_{N-1}(x)\varphi_{N-1,\rho_{N-1}(m-1)}(x)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(m-1-n^{\prime})\big)},$ (3.44)
where we have used the fact that both $W_{N-1}(x)$ and $\varphi_{N-1,\rho_{N-1}(m-1)}$ are invariant under translations $x\to x+(t,\ldots,t)$ for any $t\in{\mathbb{C}}$. Using (3.7) and (3.40) for $N\to N-1$, we easily arrive at the right-hand side of (3.40). ∎
To establish quasi-invariance, we make use of the fact that $\varphi_{N}$ is $S_{N}$-invariant in the sense of the following lemma.
Lemma 3.8.
We have
$\varphi_{N}(\sigma(x),\sigma(y))=(-)^{|\sigma|}\varphi_{N}(x,y),\ \ \ \sigma\in S_{N}.$ (3.45)
Proof.
From (3.28) and the induction assumption, we see immediately that the statement holds true whenever $\sigma\in S_{N-1}$. Hence, we need only to consider the special case $\sigma=\sigma_{N-1}$. By a direct computation, similar to the one in the proof of Proposition 3.7, we obtain
$\begin{split}\varphi_{N,\sigma_{N-1}(\rho_{N}(m-1))}(x)&=-\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{1}{\prod_{k=1}^{N}\prod_{j=1}^{\min(k,N-1)-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in\big)}\\ &\quad\cdot\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{N-1}-x_{N}+in\big)}\\ &=\varphi_{N,\rho_{N}(m-1)}(\sigma_{N-1}(x)).\end{split}$ (3.46)
(Note that only the term in (3.41) corresponding to $\nu^{\prime}=\rho_{N-1}$, $n_{1}=\cdots=n_{N-2}=m-1$ and $n_{N-1}=0$ contributes.) It follows that $\varphi_{N}(x,y)$ and $\varphi_{N}(\sigma_{N-1}(x),\sigma_{N-1}(y))$ are both (finite) series of the form
$e^{\frac{2\pi i}{a}(x-i\rho_{N}(m-1),y)}f_{N}(x,y),\ \ \ f_{N}(x,y)=\Gamma_{0}(x)+\sum_{\nu>0}\Gamma_{\nu}(x)e^{-\frac{2\pi}{a}(\nu,y)},$ (3.47)
where $\Gamma_{0}=\varphi_{N,\rho_{N}(m-1)}$ and $\nu>0$ means that $\nu=\sum_{1\leq j<k\leq N}l_{jk}(e_{j}-e_{k})$ with $l_{jk}\geq 0$ and not all equal to zero. From Proposition 3.7, we get that the functions $\varphi_{N}(\sigma(x),y)$, $\sigma\in S_{N}$, form a linearly independent set; and, since the pre-factor $\prod_{j<k}2\sinh(\pi(y_{j}-y_{k}))$ in (3.30) is antiperiodic under shifts $y_{j}\to y_{j}-i$ and the Macdonald operators $D_{N}^{r}$ are $S_{N}$-invariant, it follows that $\varphi_{N}(x,y)$ and $\varphi_{N}(\sigma_{N-1}(x),\sigma_{N-1}(y))$ satisfy the difference equation
$D_{N}(y)\varphi(x,y)=\varphi(x,y)\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}},\ \ \ D_{N}(y)\equiv D^{1}_{N}(q^{i2y};q^{2},q^{2(1-m)}).$ (3.48)
For the corresponding series $f_{N}$, this translates into the difference equation $L_{N}(y)f(x,y)=f(x,y)\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}}$ given by the difference operator
$\begin{split}L_{N}(y)&\equiv e^{-\frac{2\pi i}{a}(x-i\rho_{N}(m-1),y)}D_{N}(y)e^{\frac{2\pi i}{a}(x-i\rho_{N}(m-1),y)}\\ &=\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}}\prod_{k<j}\frac{1-e^{\frac{2\pi}{a}(y_{j}-y_{k}+i(m-1))}}{1-e^{\frac{2\pi}{a}(y_{j}-y_{k})}}\cdot\prod_{k>j}\frac{1-e^{\frac{2\pi}{a}(y_{k}-y_{j}-i(m-1))}}{1-e^{\frac{2\pi}{a}(y_{k}-y_{j})}}\cdot e^{-i\partial_{y_{j}}}.\end{split}$ (3.49)
By expanding the coefficients in power series in $e^{\frac{2\pi}{a}(y_{k}-y_{j})}$, $1\leq j<k\leq N$, this, together with the leading coefficient $\Gamma_{0}$, is readily seen to uniquely determine $f_{N}$. This essentially amounts to the uniqueness result in Theorem 2.3 in [NS12], wherein a detailed proof can be found. ∎
Proposition 3.9.
For all $1\leq j<k\leq N$ and $s=1,\ldots,m-1$, we have the vanishing property
$\varphi_{N}\left(x;y+\frac{is}{2}(e_{j}-e_{k})\right)=\varphi_{N}\left(x;y-\frac{is}{2}(e_{j}-e_{k})\right),\ \ \ e^{\frac{2\pi}{a}(y_{j}-y_{k})}=1.$ (3.50)
Proof.
Thanks to Lemma 3.8, it suffices to establish the vanishing property when $j=1$ and $k=2$; and, in this case, it follows immediately from (3.28) and our induction assumption. ∎
Comparing Props. 3.7 and 3.9 with the properties (2.16) and (2.18)–(2.19), which uniquely characterises the self-dual BA-function $\psi_{N}$, we find the following result.
Proposition 3.10.
Given $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
$\begin{split}\varphi_{N}(a,m;x,y)&=\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{\psi_{N}(\exp(-i\pi/a),m-1;ix,iy)}{\prod_{1\leq j<k\leq N}\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}.\end{split}$ (3.51)
Finally, by combining Props. 3.6 and 3.10, we arrive at the statement in Thm. 3.1 for arbitrary NN; and, by using the above expression to express $\varphi_{N}$ and $\varphi_{N-1}$ in (3.28) in terms of $\psi_{N}$ and $\psi_{N-1}$, respectively, we easily verify the claim in Prop. 3.2.
Acknowledgements
Financial support from the Swedish Research Council (Project-id 2024-05649_VR) is gratefully acknowledged.