### v2

# Baker–Akhiezer specialisation of joint eigenfunctions for hyperbolic relativistic Calogero–Moser Hamiltonians

Abstract
In earlier joint work with Ruijsenaars, we constructed and studied symmetric joint eigenfunctions $J_{N}$ for quantum Hamiltonians of the hyperbolic relativistic $N$-particle Calogero–Moser system. For generic coupling values, they are non-elementary functions that in the $N=2$ case essentially amount to a ‘relativistic’ generalisation of the conical function specialisation of the Gauss hypergeometric function $_{2}F_{1}$. In this paper, we consider a discrete set of coupling values for which the solution to the joint eigenvalue problem is known to be given by functions $\psi_{N}$ of Baker–Akhiezer type, which are elementary, but highly nontrivial, functions. Specifically, we show that $J_{N}$ essentially amounts to the antisymmetrisation of $\psi_{N}$ and, as a byproduct, we obtain a recursive construction of $\psi_{N}$ in terms of an iterated residue formula.
1 Introduction
[…]
$S_{r}(g;x)\equiv\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}f_{-}(g;x_{j}-x_{k})\prod_{l\in I}\exp(-i\hbar\beta\partial_{x_{l}})\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}f_{+}(g;x_{j}-x_{k}),$ (1.1)
where
$f_{\pm}(g;z)=\left(\frac{\sinh(\mu(z\pm i\beta g)/2)}{\sinh(\mu z/2)}\right)^{1/2},$ (1.2)
is of particular importance. Here, $r=1,\ldots,N$, it is natural to view $\beta>0$ as $1/mc$, with $m$ the particle mass and $c$ the speed of light, and in the nonrelativistic limit $c\to\infty$ pairwise commuting Hamiltonians of the nonrelativistic hyperbolic Calogero–Moser system are recovered; see e.g. the surveys [Rui99, Hal25] and references therein.
[…]
Further fundamental properties of the joint eigenfunctions were obtained by Belousov et al. in a number of recent papers [BDKK24a, BDKK24b, BDKK24c, BDKK24d]. Their results include integral equations, a reflection symmetry of the coupling constant, self-duality under interchange of geometric and spectral variables as well as orthogonality and completeness relations.
[…]
$D^{r}_{N}(z;q,t)\equiv t^{r(r-1)/2}\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{tx_{j}-x_{k}}{x_{j}-x_{k}}\prod_{l\in I}T_{q,x_{l}},$ (1.3)
where $(T_{q,x_{l}}f)(x_{1},\ldots,x_{k},\ldots,x_{N})=f(x_{1},\ldots,qx_{l},\ldots,x_{N})$. (The precise relationship can be gleaned from (2.6) and (3.9).) They act on the space of symmetric polynomials in $N$ variables, on which they are simultaneously diagonalised by the symmetric $GL_{N}$ type Macdonald polynomials; see e.g. [Mac95, Nou23, Sto21]. For parameter values of the form $t=q^{m}$, $m\in{\mathbb{Z}}$, Etingof and Styrkas [ES98] and Chalykh [Cha02] constructed and studied non-symmetric joint eigenfunctions of Baker–Akhiezer (BA) type and obtained, in particular, a generalized Weyl character formula for the Macdonald polynomials, first conjectured by Felder and Varchenko [FV97], where the BA-function replaces the exponential function; see also Chalykh and Etingof [CE13] for further related results and references.
[…]
For the coupling values $g=m\hbar$ under consideration, our main result provides an expansion of symmetric joint $S_{r}$-eigenfunctions in terms of eigenfunctions that are ‘asymptotically free’ deep inside Weyl chambers $x_{\sigma(1)}>x_{\sigma(2)}>\cdots>x_{\sigma(N)}$, $\sigma\in S_{N}$. In the case of the Macdonald operators and generic parameter values, such asymptotically free solutions, often referred to as (q-)Harisch–Chandra series, have been studied in detail by Letzter and Stokman [LS08], van Meer and Stokman [vMS10] as well as Noumi and Shiraishi [NS12]; and Stokman [Sto14] derived a corresponding ($c$-function) expansion of Cherednik’s basic hypergeometric function associated to root systems. We should also note that Bullimore et al. [BKK15] showed that the partition function of the superconformal field theory $T[U(N)]$, as introduced by Gaiotto and Witten [GW09], displays the very same recursive structure as the joint $S_{r}$-eigenfunctions from [HR14]; and, identifying integration with summing residues at all (infinitely many) poles above the real axis, they obtained an expression in terms of series of an asympototically free type. Furthermore, in the $N=2$ case, di Francesco et al. [FKKSS24] recently obtained a number of interesting results on the $S_{r}$-eigenfunctions, including a similar residue computation where the relevant series are identified as (q-)Harish–Chandra series. To establish the precise connection between these expansion results and the recursive construction in [HR14] is, I believe, an interesting problem to which I plan to return to elsewhere. In the $BC_{1}$ case, the analogous connection between integral representations and series expansions was established by van de Bult et al. [vdBRS07].
The remainder of the paper is structured as follows: In Section 2, we briefly recall definitions and key properties pertaining to joint eigenfunctions of the Ruijsenaars operators (1.1) and BA-(eigen)functions for the Macdonald operators (1.3); and, in Section 3, we give the precise formulations and proofs of our results.
2 Preliminaries
2.1 Joint eigenfunctions
[…]
$a_{+}\equiv 2\pi/\mu>0,\quad\mathrm{(imaginary~period/interaction~length)}$ (2.1)
and
$a_{-}\equiv\hbar\beta>0,\quad\mathrm{(shift~step~size/Compton~wavelength)}$ (2.2)
and replace the coupling $g$ by the parameter
$b\equiv\beta g.$ (2.3)
[…]
$W_{N}(b;x)\equiv\prod_{1\leq j<k\leq N}\prod_{\delta=+,-}\frac{G(\delta(x_{j}-x_{k})+i(a_{+}+a_{-})/2)}{G(\delta(x_{j}-x_{k})+i(a_{+}+a_{-})/2-ib)},$ (2.4)
where $G(z)\equiv G(a_{+},a_{-};z)$ is the hyperbolic Gamma function from [Rui97], which is $a_{+}\leftrightarrow a_{-}$ invariant, meromorphic in $z$ and satisfies the analytic difference equations
$\frac{G(z+ia_{\delta}/2)}{G(z-ia_{\delta}/2)}=2\cosh(\pi z/a_{-\delta}),\ \ \ \delta=+,-;$ (2.5)
see loc. cit. Prop. III.1 and Prop. III.2. (Unless needed, we suppress dependence of $G$ and functions constructed from $G$ on the parameters $a_{\pm}$.) Indeed, using these difference equations, it is readily verified that
$\begin{split}A_{r,\delta}(b;x)&\equiv W_{N}(b;x)^{-1/2}H_{r,\delta}(b;x)W_{N}(b;x)^{1/2}\\ &=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh(\pi(x_{j}-x_{k}-ib)/a_{\delta})}{\sinh(\pi(x_{j}-x_{k})/a_{\delta})}\prod_{l\in I}\exp(-ia_{-\delta}\partial_{x_{l}}),\end{split}$ (2.6)
which, in contrast to $H_{r,\delta}$, act on the space of meromorphic functions. (We recall that, for $(a_{+},a_{-},b)\in(0,\infty)^{3}$ such that $b<a_{+}+a_{-}$ and $x\in{\mathbb{R}}^{N}$, the weight function is regular and positive.)
In [HR14] functions $J_{N}(b;x,y)$ having the joint eigenfunction property
$A_{r,\delta}(b;x)J_{N}(b;x,y)=J_{N}(b;x,y)\sum_{1\leq j_{1}<\cdots<j_{r}\leq N}e^{2\pi(y_{j_{1}}+\cdots+y_{j_{r}})/a_{\delta}},$ (2.7)
are constructed recursively by an explicit integral formula, with integrand built from $J_{N-1}$, the weight function $W_{N-1}$ as well as the kernel function
${\cal S}_{N}^{\sharp}(b;x,y)\equiv\prod_{j=1}^{N}\prod_{k=1}^{N-1}\frac{G(x_{j}-y_{k}-ib/2)}{G(x_{j}-y_{k}+ib/2)}.$ (2.8)
More precisely, with $J_{1}(x,y)\equiv\exp(2\pi ixy/a_{+}a_{-})$ as the starting point for the recursion, $J_{N}$, $N>1$, is given by
$J_{N}(b;x,y)=\frac{\exp\left(\frac{2\pi i}{a_{+}a_{-}}y_{N}\sum_{j=1}^{N}x_{j}\right)}{(N-1)!}\\ \cdot\int_{{\mathbb{R}}^{N-1}}dzW_{N-1}(b;z){\cal S}^{\sharp}_{N}(b;x,z)J_{N-1}(b;z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$ (2.9)
From Prop. III.2 in [Rui97], we recall that the hyperbolic gamma function is scale invariant, in the sense that
$G(\lambda a_{+},\lambda a_{-};\lambda z)=G(a_{+},a_{-};z),\quad\lambda\in(0,\infty).$ (2.10)
Since $J_{N}$ is constructed almost entirely from $G(z)$, it is easily seen to have the invariance properties
$J_{N}(a_{+},a_{-},b;x,y)=J_{N}(a_{-},a_{+},b;x,y),$ (2.11)
$J_{N}(\lambda a_{+},\lambda a_{-},\lambda b;\lambda x,\lambda y)=\lambda^{N(N-1)/2}J_{N}(a_{+},a_{-},b;x,y).$ (2.12)
Hence, we may and shall restrict attention to
$a_{+}=1,\ \ \ a_{-}\equiv a\geq 1,$ (2.13)
without any loss of generality. Moreover, to facilitate the comparison with the pertinent BA-function, it is expedient to renormalize $J_{N}$ and introduce the function
$\Phi_{N}(a,b;x,y)\equiv\left(\frac{G(1,a;ib-i(1+a)/2)}{\sqrt{a}}\right)^{N-1}J_{N}(1,a,b;x,y),$ (2.14)
which, in particular, satisfies the simple duality relation
$\Phi_{N}(a,b;x,y)=\Phi_{N}(a,1+a-b;y,x),$ (2.15)
as conjectured in [HR14], verified for $N\leq 3$ in [HR18a] and proven for arbitrary $N$ in [BDKK24b] (see Thm. 5 and Eqs. (1.53)–(1.54)).
2.2 Baker–Akhiezer functions
Take $q\in\mathbb{C}^{\times}\equiv{\mathbb{C}}\setminus\{0\}$, $m\in\mathbb{Z}_{+}$ and let $\psi_{N}$ be a function of $x,y\in\mathbb{C}^{N}$ that is of the form
$\psi_{N}(x,y)=q^{2(x,y)}\sum_{\nu}\psi_{N,\nu}(x)q^{2(\nu,y)},$ (2.16)
where the sum extends over weight vectors
$\nu=\sum_{1\leq j<k\leq N}\left(\frac{m}{2}-l_{jk}\right)(e_{j}-e_{k}),\ \ \ l_{jk}=0,\ldots,m,$ (2.17)
with $\{e_{1},\ldots,e_{N}\}$ the standard ON-basis in $\mathbb{R}^{N}$. Suppose, in addition, that $\psi_{N}$ satisfies the vanishing condition
$\psi_{N}(x,y+s(e_{j}-e_{k})/2)-\psi_{N}(x,y-s(e_{j}-e_{k})/2)=0\ \ \text{when}\ \ q^{2(y_{j}-y_{k})}=1,$ (2.18)
for all $1\leq j<k\leq N$ and $s=1,\ldots,m$. Then it is called a Baker–Akhiezer (BA) function associated with the root system $A_{N-1}$ and (positive) integer parameter $m$. Existence is established by different constructions in Sect. 5 of [ES98] and Sect. 3.2 of [Cha02]. We note that the iterated residue formula obtained in this paper yields an additional (constructive) existence proof.
Assuming $q$ is not a root of unity, we know from [ES98, Cha02] that a function $\psi_{N}$ with the above properties is unique up to a choice of normalisation; and that, imposing the normalisation condition
$\psi_{N,\rho_{N}}(x)=\prod_{1\leq j<k\leq N}\prod_{n=1}^{m}[n+x_{k}-x_{j}],\ \ \ [z]\equiv q^{z}-q^{-z},$ (2.19)
yields a self-dual BA function, in the sense that $\psi_{N}(x,y)=\psi_{N}(y,x)$. Here, we have used the standard notation
$\rho_{N}\equiv\rho_{N}(m)=\frac{m}{2}\sum_{1\leq j<k\leq N}(e_{j}-e_{k})=\frac{m}{2}\sum_{j=1}^{N}(N-2j+1)e_{j},$ (2.20)
obtained by setting all $l_{ij}=0$ in (2.17). The self-dual BA-function satisfies the bispectral system of $q$-difference equations
$D^{r}_{N}(q^{2y};q^{2},q^{-2m})\psi(x,y)$ $=\psi(x,y)\sum_{1\leq j_{1}<\cdots<j_{r}\leq N}q^{2(x_{j_{1}}+\cdots+x_{j_{r}})},$ (2.21)
$D^{r}_{N}(q^{2x};q^{2},q^{-2m})\psi(x,y)$ $=\psi(x,y)\sum_{1\leq j_{1}<\cdots<j_{r}\leq N}q^{2(y_{j_{1}}+\cdots+y_{j_{r}})},$ (2.22)
with $r=1,\ldots,N$, $q^{2x}\equiv(q^{2x_{1}},\ldots,q^{2x_{N}})$ and $q^{2y}\equiv(q^{2y_{1}},\ldots,q^{2y_{N}})$.
3 Baker–Akhiezer specialisation
When $b=m\in\mathbb{Z}_{+}$, the weight function $W_{N}$ (2.4) and kernel function ${\cal S}^{\sharp}_{N}$ (2.8) reduce to elementary functions. More precisely, from the difference equation (2.5) satisfied by $G(z)$, we see that
$W_{N}(1,a,m;x)=\prod_{1\leq j<k\leq N}\sinh\left(\frac{\pi}{a}(x_{k}-x_{j})\right)\prod_{n=-m+1}^{m-1}2\sinh\left(\frac{\pi}{a}(x_{k}-x_{j}+in)\right)$ (3.1)
and
${\cal S}^{\sharp}_{N}(1,a,m;x,y)=\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}\frac{1}{2\cosh\big(\frac{\pi}{a}(x_{j}-y_{k}+i(m-2n-1)/2)\big)}.$ (3.2)
Up to an overall numerical factor, $W_{N}(1,a,m;x)$ equals the specialisation $\Delta(q^{2ix};q^{2},q^{2m})$ of Macdonald’s weight function $\Delta$, while ${\cal S}^{\sharp}_{N}(1,a,m;x,y)$ is obtained by a similar specialisation of his product function $\Pi$ after removing a factor depending only on $\sum_{j=1}^{N}(x_{j}-y_{j})$; cf. Eqs. (9.2) and (2.5), respectively, in Section VI of [Mac95]. Furthermore, (2.5) and the formula $G(1,a;i(1-a)/2)=1/\sqrt{a}$ (see Eq. (3.38) in [Rui97]) entail that
$\frac{G(1,a;im-i(1+a)/2)}{\sqrt{a}}=\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}.$ (3.3)
Consequently, the $b=m$ specialisation of $\Phi_{N}$ is given by recursively by
$\Phi_{N}(m;x,y)=\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a(N-1)!}e^{\frac{2\pi i}{a}y_{N}(x_{1}+\cdots+x_{N})}\\ \cdot\int_{\mathbb{R}^{N-1}}dz\prod_{1\leq j<k\leq N-1}\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)\prod_{n=-m+1}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j}+in)\right)\\ \cdot\frac{\Phi_{N-1}(m;z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\cosh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)}$ (3.4)
with $\Phi_{1}(x,y)=\exp\left(\frac{2\pi}{a}xy\right)$. We note that this is effectively a basic hypergeometric integral, since the weight and kernel functions can be expressed in terms of the product
$\prod_{n=0}^{m-1}(q^{n+z}-q^{-n-z})=(-1)^{m}q^{-m(m-1)/2-mz}\frac{(q^{2z};q^{2})_{\infty}}{(q^{2(m+z)};q^{2})_{\infty}},$ (3.5)
with
$q=e^{-i\pi/a},\ \ \ a\in\mathbb{R}.$ (3.6)
To be precise, while the (modified) $q$-Gamma function $(q^{2z};q^{2})_{\infty}$ is only well-defined for $|q|<1$, the above ratio clearly extends to $|q|=1$. For accounts of (multivariable) basic hypergeometric integrals, see e.g. the overview by Koornwinder and Stokman [KS21] and the survey by Schlosser [Sch21].
Upon setting $(a_{+},a_{-},b)=(1,a,m)$, the $a$-antiperiodicity of the functions involved entails that the eigenvalue equations (2.7), which are are also satisfied by $\Phi_{N}$, become trivial for $\delta=+$. As we now show, the remaining equations, corresponding to $\delta=$-, are closely related to the Macdonald eigenvalue equations (2.21), of which the BA-functions are solutions.
Letting
$\mathcal{\delta}_{N}(x;m)=\prod_{1\leq j<k\leq N}\prod_{n=-m}^{m}[x_{j}-x_{k}+n],$ (3.7)
we recall the well-known similarity transform
$\mathcal{\delta}_{N}(x;m-1)^{-1}D^{r}_{N}(q^{2x};q^{2},q^{-2(m-1)})\mathcal{\delta}_{N}(x;m-1)=D^{r}_{N}(q^{2x};q^{2},q^{2m}),$ (3.8)
which is easily verified by a direct computation; and with the parametrisation (3.6) in effect it is readily seen that
$q^{mr(1-N)}D^{r}_{N}(q^{2ix};q^{2},q^{2m})=\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=r\end{subarray}}\prod_{\begin{subarray}{c}j\in I\\ k\notin I\end{subarray}}\frac{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}-im)\big)}{\sinh\big(\frac{\pi}{a}(x_{j}-x_{k})\big)}\prod_{l\in I}\exp(-i\partial_{x_{l}}).$ (3.9)
[…]
$\mathcal{\delta}_{N}(i\sigma(x);m-1)=(-)^{\sigma}\mathcal{\delta}_{N}(ix;m-1),\ \ \ \sigma\in S_{N},$ (3.10)
which is clear from (3.7). Moreover, given that $\Phi_{N}(a,m;x,y)$ is manifestly symmetric in $x$ (cf. (2.9) and (2.14)), the coefficients in the linear combination should be of the form $c(y)(-)^{\sigma}$, with $c$ independent of $\sigma$. In this section, we substantiate this expectation by proving the following theorem.
Theorem 3.1.
Let $m\in\mathbb{Z}_{+}$ and let a>m1a>m-1 be such that $\exp(-i\pi/a)$ is not a root of unity. Then, we have
$\Phi_{N}(a,m;x,y)\\ =C_{N}(a,m)\frac{\sum_{\sigma\in S_{N}}(-)^{\sigma}\psi_{N}(e^{-i\pi/a},m-1;i\sigma(x),iy)}{\prod_{1\leq j<k\leq N}2\sinh\big(\frac{\pi}{a}(y_{k}-y_{j})\big)\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}$ (3.11)
with the constant
$C_{N}(a,m)=(i^{m})^{N(N-1)/2}\left(\frac{a}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{(N-1)(N-2)/2}.$ (3.12)
The first main observation in our proof of Thm. 3.1, as detailed below, is that the product function $\Phi_{N}(a,m;x,y)\prod_{1\leq j<k\leq N}2\sinh(\pi(y_{j}-y_{k}))$ can be rewritten as the symmetrisation of a function $\varphi_{N}(a,m;x,y)$ that is given recursively by an iterated residue formula.
More specifically, the formula involves denominator factors $\sinh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)$, producing poles at
$z_{k}=x_{j}+i(m-2n-1)/2+i\ell a,\ \ \ n=0,\ldots,m-1,\ \ \ell\in\mathbb{Z},$ (3.13)
where $j=1,\ldots,N$ and $k=1,\ldots,N-1$. Assuming that ${\rm Re}\,(x_{j}-x_{k})\neq 0$ for all $1\leq j<k\leq N$ and a>m1a>m-1, we can find contours $\gamma_{k}$ that only encircle the poles corresponding to a fixed value of $k=1,\ldots,N-1$ and $\ell=0$; see the discussions preceding Props. 3.3 and 3.6 for explicit examples. Up to normalisation, the recursive formula defining $\varphi_{N}$ can then be obtained from (3.4) by the substitutions $\Phi\to\varphi$, ${\mathbb{R}}^{N-1}\to\gamma_{1}\times\cdots\times\gamma_{N-1}$ and $\cosh\to\sinh$.
The second main observation, which more or less completes the proof of Thm. 3.1, is that $\varphi_{N}$ essentially amounts to a renormalisation of $\psi_{N}$ and, as a byproduct, we thus find the following recursive construction of the self-dual BA-function $\psi_{N}$.
Proposition 3.2.
Under the above assumptions on $m$ and $a$, the BA-function $\psi_{N}$ is obtained from $\psi_{N-1}$ by the iterated residue formula
$\psi_{N}(e^{-i\pi/a},m;ix,iy)\\ =\left(\frac{\prod_{n=1}^{m}\sin(\pi n/a)}{a(-i)^{m+1}}\right)^{N-1}\frac{W_{N}(1,a,m+1;z)}{\prod_{1\leq j<k\leq N}2\sinh\left(\frac{\pi}{a}(x_{k}-x_{j})\right)}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\frac{\psi_{N-1}(e^{-i\pi/a},m;iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j<k\leq N-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n)\big)\big)},$ (3.14)
where $\underline{\gamma}\equiv\gamma_{1}\times\cdots\times\gamma_{N-1}$.
This result provides a natural generalisation of the iterated residue formulae for BA-functions for rational and trigonometric Calogero–Moser–Sutherland operators obtained by Felder and Veselov [FV09].
In the two subsections below, we provide our proofs of Thm. 3.1 and Prop. 3.2. To begin with, we establish the $N=2$ cases of the results. These we then use as the basis for proofs by induction on $N$. To shorten some of the formulae involved, we typically suppress dependence on the parameters $a$ and $m$.
3.1 The $N=2$ case
Setting $N=2$ in (3.4), we deduce
$\frac{2a\sinh\big(\frac{\pi}{a}(y_{1}-y_{2})\big)}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\Phi_{2}(x,y)\\ =\int_{\mathbb{R}}dz\frac{e^{\frac{2\pi i}{a}(y_{1}-y_{2})(z-ia/2)}-e^{\frac{2\pi i}{a}(y_{1}-y_{2})(z+ia/2)}}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\cosh\big(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2)\big)}\\ =(-1)^{m}\left(\int_{{\mathbb{R}}-ia/2}dz-\int_{{\mathbb{R}}+ia/2}dz\right)\frac{e^{\frac{2\pi i}{a}(y_{1}-y_{2})z}}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2)\big)}.$ (3.15)
We note that a>m1a>m-1 and $x,y\in\mathbb{R}^{2}$ ensures that $\Phi_{2}$ is well-defined and that the only poles of the integrand located within the strip $|{\rm Im}\,z|<a/2$ are
$z=x_{j}+i(m-2n-1)/2,\ \ \ j=1,2,\ \ n=0,\ldots,m-1.$ (3.16)
At first, we assume $x=(x_{1},x_{2})\in{\mathbb{R}}^{2}$ is such that $x_{1}\neq x_{2}$, which ensures that we can introduce a contour $\gamma_{1}$ encircling only the poles (3.16) with $j=1$ counterclockwise. For example, with $\delta\equiv|x_{1}-x_{2}|>0$, we can take the contour consisting of line segments $x_{1}+u\pm ia/2$, $|u|\leq\delta/2$, and $x_{1}\pm\delta/2+iv$, $|v|\leq a/2$. Introducing the function
$\varphi_{2}(x,y)\equiv e^{\frac{2\pi i}{a}y_{2}(x_{1}+x_{2})}\int_{\gamma_{1}}dz\frac{e^{\frac{2\pi i}{a}(y_{1}-y_{2})z}}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2)\big)},$ (3.17)
the following result is an easy consequence of (3.15) and Cauchy’s residue theorem.
Proposition 3.3.
For $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
$\Phi_{2}(x,y)=(-1)^{m}\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\frac{\sum_{\sigma\in S_{2}}\varphi_{2}(\sigma(x),y)}{2\sinh\big(\frac{\pi}{a}(y_{1}-y_{2})\big)}.$ (3.18)
We proceed to establish the precise connection between $\varphi_{2}$ and the self-dual BA-function $\psi_{2}$. As detailed in the following proposition, when applying Cauchy’s residue thm. to (3.17) we obtain a series expansion for $\varphi_{2}$ analogous to the expansion of $\psi_{2}$ given by the $N=2$ instance of (2.16).
Proposition 3.4.
Assuming that $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
$\varphi_{2}(x,y)=e^{\frac{2\pi i}{a}(x,y)}\sum_{l=0}^{m-1}\varphi_{2,l}(x)e^{\frac{2\pi}{a}\left(\frac{m-1}{2}-l\right)(y_{1}-y_{2})}$ (3.19)
with
$\varphi_{2,l}(x)=\frac{a(-i)^{m}}{\prod_{n\neq l}2\sin\big(\frac{\pi}{a}(n-l)\big)}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))\big)}.$ (3.20)
We note that the above series essentially amounts to a special (terminating) case of the basic hypergeometric series
$_{2}\phi_{1}\left(\,\begin{matrix}a,\,b\\ \,c\end{matrix}\,\Big|\,q,z\right)=\sum_{l=0}^{\infty}\frac{(a:q)_{l}(b:q)_{l}}{(c;q)_{l}(q;q)_{l}}z^{l},$ (3.21)
where the $q$-shifted factorial is given by $(a;q)_{0}=1$ and $(a;q)_{l}=(1-a)\cdots(1-aq^{l-1})$ when $l>0$. To verify this claims, it suffices to observe that
$\begin{split}\frac{1}{\prod_{n\neq l}2\sin\left(\frac{\pi}{a}(n-l)\right)}&=\frac{1}{\prod_{n=1}^{m-1}2\sin\left(\frac{\pi}{a}n\right)}\prod_{n=0}^{l-1}\frac{2\sin\left(\frac{\pi}{a}(m-1-n)\right)}{2\sin\left(\frac{\pi}{a}(-n-1)\right)}\\ &=\frac{t_{m}^{l}}{\prod_{n=1}^{m-1}2\sin\left(\frac{\pi}{a}n\right)}\frac{(q^{2}/t_{m}^{2};q^{2})_{l}}{(q^{2};q^{2})_{l}}\end{split}$ (3.22)
and similarly that
$\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))\big)}\\ =t_{m}^{l}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}\frac{(q^{2}X_{2}/t_{m}^{2}X_{1};q^{2})_{l}}{(q^{2}X_{2}/X_{1};q^{2})_{l}},$ (3.23)
where we have used the parametrisation (3.6) for $q$ and introduced
$t_{m}\equiv e^{-\frac{i\pi}{a}m}=q^{m}$ (3.24)
as well as
$X_{j}\equiv e^{\frac{2\pi}{a}x_{j}}=q^{i2x_{j}},\ \ Y_{j}\equiv e^{\frac{2\pi}{a}y_{j}}=q^{i2y_{j}},\ \ \ j=1,2.$ (3.25)
Indeed, it is clear from the above computations that
$\begin{split}\varphi_{2}(x,y)&=\frac{a(-i)^{m}\exp\left(\frac{2\pi i}{a}(x-i\rho_{2}(m-1),y)\right)}{\prod_{n=1}^{m-1}2\sin(\pi n/a)\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}\\ &\quad\cdot{}_{2}\phi_{1}\left(\,\begin{matrix}q^{2}/t_{m}^{2},\,q^{2}X_{2}/t_{m}^{2}X_{1}\\ \,q^{2}X_{2}/X_{1}\end{matrix}\,\Big|\,q^{2},t_{m}^{2}Y_{2}/Y_{1}\right).\end{split}$ (3.26)
Starting from the $q$-difference eigenvalue equation satisfied by $_{2}\phi_{1}$, it would now be a simple exercise to show that $\varphi_{2}$ is an eigenfunction of the $N=2$ Macdonald operator $D_{2}^{1}$ for suitable parameter values. One could also invoke the duality relation (2.15) and linear independence of $\varphi_{2}(\sigma(x),y)$, $\sigma\in S_{2}$, for generic $x\in\mathbb{R}^{2}$; see Lemma 3.8 below. Either way, one could then use a uniqueness result such as Prop. 4.1 in [Nou23] for ‘asymptotically free’ eigenfunctions of $D_{2}^{1}$ to arrive at the connection between $\varphi_{2}$ and $\psi_{2}$. We note that such uniqueness results for arbitrary $N$ can be found in the paper of Letzter and Stokman [LS08] (in a general root system setting) as well as that of Noumi and Shiraishi [NS12].
Here, we rely instead on the defining properties of $\psi_{2}$ and show that $\varphi_{2}$ has vanishing properties similar to those given by (2.18) with $N=2$.
Proposition 3.5.
For $s=1,\ldots,m-1$, we have
$\varphi_{2}\left(x,y+is(e_{1}-e_{2})/2\right)-\varphi_{2}\left(x,y-is(e_{1}-e_{2})/2\right)=0\ \ \text{when}\ \ e^{\frac{2\pi}{a}(y_{1}-y_{2})}=1.$ (3.27)
Proof.
Up to an overall factor $2\exp(\frac{2\pi i}{a}(y_{2}(x_{1}+x_{2})))$, the LHS of (3.27) is given by the integral
$\int_{\gamma_{1}}dz\frac{\exp\big(\frac{2\pi i}{a}(y_{1}-y_{2})z\big)\sinh(\frac{\pi s}{a}(x_{1}+x_{2}-2z))}{\prod_{j=1}^{2}\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{j}-z+i(m-2n-1)/2))},$ (3.28)
which, by Cauchy’s theorem, equals
$(-i)^{m}e^{\frac{2\pi i}{a}(y_{1}-y_{2})x_{1}}\sum_{l=0}^{m-1}\frac{a}{\prod_{n\neq l}2\sin(\frac{\pi}{a}(n-l))}\\ \cdot\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}-i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}e^{\frac{\pi}{a}(m-2l-1)(y_{1}-y_{2})}.$ (3.29)
Since $e^{-\frac{\pi}{a}(m-2l-1)(y_{1}-y_{2})}=e^{\frac{\pi}{a}(m-1)(y_{1}-y_{2})}$ when $e^{\frac{2\pi}{a}(y_{1}-y_{2})}=1$, we only have to show that
$\sum_{l=0}^{m-1}\frac{a}{\prod_{n\neq l}2\sin(\frac{\pi}{a}(n-l))}\frac{\sinh(\frac{\pi s}{a}(x_{2}-x_{1}+i(m-2l-1))}{\prod_{n=0}^{m-1}2\sinh(\frac{\pi}{a}(x_{2}-x_{1}+i(n-l))}=0.$ (3.30)
To this end, we note that $s<m$ entails that the left-hand side is a meromorphic function in $x_{1}-x_{2}$ that is bounded away from its poles and decays exponentially as $|{\rm Re}\,(x_{1}-x_{2})|\to\infty$. Hence, Liouville’s theorem will imply the above identity once we show that the residues at the (simple) poles
$x_{2}-x_{1}+i\ell=i\ell^{\prime}a,\ \ \ \ell=-m+1,\ldots,m-1,\ \ \ell^{\prime}\in\mathbb{Z},$ (3.31)
all vanish; and, by iaia-(anti)periodicity, we may and shall restrict attention to $\ell^{\prime}=0$. For a fixed $\ell=-m+1,\ldots,m-1$, we observe that there exists $n=0,\ldots,m-1$ such that $n-l=\ell$ if and only if
$l=\max(-\ell,0),\ldots,\min(m-1,m-1-\ell).$ (3.32)
It follows that the residue of the left-hand side of (3.30) at $x_{2}-x_{1}+i\ell=0$ is proportional to
$\sum_{l=\max(-\ell,0)}^{\min(m-1,m-1-\ell)}f_{l},\ \ \ f_{l}(a)\equiv\frac{2a}{\prod_{n\neq l}\sin(\frac{\pi}{a}(n-l))}\frac{\sin(\frac{\pi s}{a}(2l+\ell-m+1))}{\prod_{n\neq l+\ell}\sin(\frac{\pi}{a}(n-l-\ell))}.$ (3.33)
On the set of integers (3.32), we have the involution $\sigma:l\mapsto m-1-\ell-l$. Since $f_{\sigma(l)}=-f_{l}$, as is easily checked, this clearly implies that the residue vanishes. ∎
Setting $l=0$ in (3.20), we find that
$\varphi_{2,0}(x)=\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}.$ (3.34)
Comparing (3.34), (3.27) and (3.26) with the $N=2$ instances of (2.19), (2.18) and (2.16), respectively, the uniqueness of the BA-function, once a normalisation has been fixed, implies that
$\varphi_{2}(x,y)=\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\frac{\psi_{2}(\exp(-i\pi/a),m-1;i\sigma(x),iy)}{\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{2}-x_{1}+in)\big)}.$ (3.35)
When combined with Prop. 3.3, we obtain the $N=2$ case of Thm. 3.1; and a comparison with (3.17) yields the $N=2$ instance of Prop. 3.2.
3.2 The inductive step
We can now use the $N=2$ case as the basis for proofs by induction. More precisely, we consider $N\geq 3$ and assume that Thm. 3.1 along with Props. 3.63.7, Lemma 3.8 and Props. 3.93.10 hold true when $N$ is replaced by N1N-1. (In the $N=2$ case, the pertinent results have all been established above.)
To begin with, we require that $x=(x_{1},\ldots,x_{N})\in{\mathbb{R}}^{N}$ with $x_{j}\neq x_{k}$ for all $1\leq j<k\leq N$. Letting $\delta=\min\{|x_{j}-x_{k}|\mid 1\leq j<k\leq N\}$, we construct $N$ contours $\gamma_{j}$, $j=1,\ldots,N$, from the line segments $x_{j}+u\pm ia/2$, $|u|\leq\delta/2$, and $x_{j}\pm\delta/2+iv$, $|v|\leq a/2$; and define the function
$\varphi_{N}(x,y)\equiv e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\prod_{1\leq j<k\leq N-1}\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)\prod_{n=-m+1}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j}+in)\right)\\ \cdot\frac{\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.36)
with $\underline{\gamma}\equiv\gamma_{1}\times\cdots\times\gamma_{N-1}$, and where the $z_{j}$-contour $\gamma_{j}$ only encircles the simple poles
$z_{j}=x_{j}+i(m-2n-1)/2,\ \ \ n=0,\ldots,m-1.$ (3.37)
The arbitrary-$N$ generalisation of Prop. 3.3 now follows.
Proposition 3.6.
Assuming $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we get
$\prod_{1\leq j<k\leq N}2\sinh(\pi(y_{j}-y_{k}))\cdot\Phi_{N}(x,y)\\ =(-1)^{mN(N-1)/2}\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}\sum_{\sigma\in S_{N}}\varphi_{N}(\sigma(x),y).$ (3.38)
Proof.
To ease the notation, we suppress dependence on the parameters $a$ and $m$ throughout the proof. Substituting (3.4) in the left-hand side of (3.38), using (3.1)–(3.2) to reduce the size of the resulting expression and invoking the above statement after taking $N\to N-1$, we obtain
$\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}\prod_{j=1}^{N-1}2\sinh(\pi(y_{j}-y_{N}))\frac{e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}}{(N-1)!}\\ \cdot\sum_{\sigma\in S_{N-1}}\int_{\mathbb{R}^{N-1}}dzW_{N-1}(z){\cal S}^{\sharp}_{N}(x,z)\varphi_{N-1}(\sigma(z),(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$
Thanks to the $S_{N-1}$-invariance of the domain of integration $\mathbb{R}^{N-1}$ as well as $W_{N-1}(z)$ and ${\cal S}^{\sharp}_{N}(x,z)$, we can replace the sum over $\sigma\in S_{N-1}$ by a factor (N1)!(N-1)!. By also substituting $2\sinh(\pi(y_{j}-y_{N}))=e^{\pi(y_{j}-y_{N})}-e^{-\pi(y_{j}-y_{N})}$, we thus obtain
$\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\sum_{\delta\in\{\pm 1\}^{N-1}}\delta_{1}\cdots\delta_{N-1}\\ \cdot\int_{\mathbb{R}^{N-1}}dzW_{N-1}(z){\cal S}^{\sharp}_{N}(x,z)e^{\pi\sum_{j=1}^{N-1}\delta_{j}(y_{j}-y_{N})}\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N})).$ (3.39)
To rewrite the integral further we shall make use of the invariance property
$e^{\pi\sum_{j=1}^{N-1}\delta_{j}u_{j}}\varphi_{N-1}(z,u)=\varphi_{N-1}(z-(ia/2)\delta,u)\prod_{j=1}^{N-1}\delta_{j}^{m(N-2)}.$ (3.40)
When N1=2N-1=2 it is a simple consequence of Prop. 3.4 and, using this as the basis for an induction argument, the validity of the formula in the general-$N$ case is readily inferred from (3.49)–(3.50) and the observation that $W_{N-1}(z+(ia/2)\delta)=W_{N-1}(z)$. Substituting (3.40) (with $u_{j}=y_{j}-y_{N}$) in the above integral, taking $z\to z+(ia/2)\delta$, under which $W_{N-1}(z)$ is invariant, and noting that
${\cal S}^{\sharp}_{N}(x,z+(ia/2)\delta)=\frac{(-1)^{mN(N-1)/2}\prod_{j=1}^{N-1}\delta_{j}^{mN}}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)},$ (3.41)
we arrive at
$(-1)^{mN(N-1)/2}\left(\frac{\prod_{n=1}^{m-1}2\sin(\pi n/a)}{a}\right)^{N-1}e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\sum_{\delta\in\{\pm 1\}^{N-1}}\delta_{1}\cdots\delta_{N-1}\\ \cdot\int_{{\mathbb{R}}-\frac{ia}{2}\delta_{1}}dz_{1}\cdots\int_{{\mathbb{R}}-\frac{ia}{2}\delta_{N-1}}dz_{N-1}\frac{W_{N-1}(z)\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.42)
where the sum of integrals amounts to
$\int_{{\cal C}^{N-1}}dz\frac{W_{N-1}(z)\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)}$ (3.43)
with $\cal C$} the contour consisting of the lines ${\mathbb{R}}\pm ia/2$, traversed from right/left to left/right. By Cauchy’s residue theorem, this one integral equals
$\sum_{\sigma\in S_{N}}\int_{\gamma_{\sigma(1)}}dz_{1}\cdots\int_{\gamma_{\sigma(N-1)}}dz_{N-1}\\ \cdot\frac{W_{N-1}(z)\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)}+R(x,y)$
for some remainder term $R(x,y)$, resulting from residues of the integrand at points $z=(z_{1},\ldots,z_{N-1})$ given by
$z_{j}=x_{i_{j}}+i(m-2n_{j}+1)/2,\ \ \ j=1,\ldots,N-1,$ (3.44)
where at least two $i_{j}$ coincide.
Referring back to (3.36), we find that it remains to prove that $R=0$. To this end, we take $N\to N-1$ in Prop. 3.10 and use the resulting formula to rewrite the integrand as
$C_{N-1}\frac{\psi_{N-1}(e^{-i\pi/a},m-1;iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j<k\leq N-1}\sinh\big(\frac{\pi}{a}(z_{j}-z_{k})\big)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.45)
where $C_{N-1}$ is a constant (for fixed $m$ and $a$), whose value is of no consequence for the arguments below. From Prop. 4.4 in [Cha02] (part (iii)), we know that $\psi_{N-1}(u,v)$ is an entire function. The presence of the factors $\sinh\big(\frac{\pi}{a}(z_{j}-z_{k})\big)$ thus entails that (iterated) residues at points (3.44) such that $i_{j}=i_{k}$ and $n_{j}=n_{k}$ for some $1\leq j\neq k\leq N-1$ all vanish. In the remaining cases we have $n_{j}\neq n_{k}$ while $i_{j}=i_{k}$. Thanks to the $S_{N-1}$-invariance of the integrand (cf. Lemma 3.8), we may and shall restrict attention to $j=1$ and $k=2$. Letting $\widetilde{\varphi}_{N-1}(z)=\psi_{N-1}(e^{-i\pi/a},m-1;iz,i(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))$, $t=x_{i_{1}}=x_{i_{2}}$ and $s=n_{1}-n_{2}$, we observe that
$\tilde{\varphi}_{N-1}\big((t+i(m-2n_{1}-1)/2,t+i(m-2n_{2}-1)/2,\ldots)\big)\\ =\tilde{\varphi}_{N-1}\big((t+i(m-n_{1}-n_{2}-1)/2-is/2,t+i(m-n_{1}-n_{2}-1)/2+is/2,\ldots)\big).$
Note that the self-duality of $\psi_{N-1}$ entails that $\widetilde{\varphi}_{N-1}(z)$ satisfies the vanishing conditions in Prop. 3.9 with $N\to N-1$ in $z$. Since each $n_{j}=0,\ldots,m-1$, we have $s=-m+1,\ldots m-1$. From the vanishing conditions for $\tilde{\varphi}$, we can thus infer that its specialisation at (3.44) with $i_{1}=i_{2}$ is invariant under the interchange $n_{1}\leftrightarrow n_{2}$. Moreover, we have
$\sinh(\pi(z_{1}-z_{2})/a)=\sinh(i\pi(n_{2}-n_{1})/a),$ (3.46)
so that the same specialisation of $\varphi(z)W_{N-1}(z)$ is antisymmetric under $n_{1}\leftrightarrow n_{2}$. In fact, more generally, when $i_{1}=\cdots=i_{M}$, $M\geq 2$, coincide, we can, similarly, show that antisymmetry extends to any permutation of the corresponding $M$ distinct integers $n_{1},\ldots,n_{M}$. (Essentially, we need only to decompose such a permutation in terms of (elementary) transpositions and appeal to the above discussion.) Either way, we see that the residues corresponding to a fixed choice of indices $i_{j}$, at least two of which coincide, and integers $n_{j}$, fixed up to permutations among the ones corresponding to equal $i_{j}$, cancel (in pairs) and consequently that $R=0$. ∎
We proceed to show that $\varphi_{N}$ (3.36) essentially amounts to the BA-function reviewed above. This requires that we establish two things: the relevant series expansion and vanishing properties.
Proposition 3.7.
We have
$\varphi_{N}(x;y)=e^{\frac{2\pi i}{a}(x,y)}\sum_{\nu}\varphi_{N,\nu}(x)e^{\frac{2\pi}{a}(\nu,y)}$ (3.47)
with the sum running over weight vectors of the form (2.17) with $m\to m-1$ and where
$\begin{split}\varphi_{N,\rho_{N}(m-1)}(x)&=\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{1}{\prod_{1\leq j<k\leq N}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}.\end{split}$ (3.48)
Proof.
By our induction assumption, we may invoke the proposition after taking $N\to N-1$. Substituting the resulting expansion (3.47) in (3.36) and making use of Cauchy’s residue theorem, we deduce
$\varphi_{N}(x;y)=e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\sum_{\nu^{\prime}}e^{\frac{2\pi}{a}(\nu^{\prime},y)}\\ \cdot\int_{\underline{\gamma}}dz\frac{W_{N-1}(z)\varphi_{N-1,\nu^{\prime}}(z)e^{\frac{2\pi i}{a}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-z_{k}+i(m-2n-1)/2)\big)}\\ =e^{\frac{2\pi i}{a}(x,y)}\sum_{\nu^{\prime}}e^{\frac{2\pi}{a}(\nu^{\prime},(y_{1},\ldots,y_{N-1}))}\sum_{n_{1}=0}^{m-1}\cdots\sum_{n_{N-1}=0}^{m-1}C_{\nu^{\prime};n_{1},\ldots,n_{N-1}}(x)e^{\frac{2\pi}{a}\sum_{k=1}^{N-1}(n_{k}-(m-1)/2)(y_{k}-y_{N})},$ (3.49)
with $\nu^{\prime}$ running over the set of weight vectors obtained after taking $N\to N-1$ and $m\to m-1$ in (2.17), and where
$C_{v^{\prime};n_{1},\ldots,n_{N-1}}(x)\\ =(-ia)^{N-1}\frac{W_{N-1}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{n^{\prime}\neq n_{k}}2\sinh\big(\frac{i\pi}{a}(n_{k}-n^{\prime})\big)}\\ \cdot\frac{\varphi_{N-1,\nu^{\prime}}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(n_{k}-n^{\prime})\big)}$ (3.50)
and we have used that $(\nu^{\prime},(y_{N},\ldots,y_{N}))=0$, since $\nu^{\prime}$ is a linear combination of vectors $e_{j}-e_{k}$ with $1\leq j<k\leq N-1$. Setting $l_{j,N-1}=m-1-n_{j}=0,\ldots,m-1$, we get the equality
$\sum_{1\leq j<k\leq N-1}\left(\frac{m-1}{2}-l_{jk}\right)(e_{j}-e_{k})+\sum_{j=1}^{N-1}\left(n_{j}-\frac{m-1}{2}\right)(e_{j}-e_{N})\\ =\sum_{1\leq j<k\leq N}\left(\frac{m-1}{2}-l_{jk}\right)(e_{j}-e_{k})$ (3.51)
and the validity of (3.47) with $\nu$ from (2.17) clearly follows.
Since the only term in (3.49) that contributes to $\varphi_{N,\rho_{N}(m-1)}(x)$ corresponds to $\nu^{\prime}=\rho_{N-1}(m-1)$ and $n_{1}=\cdots=n_{N-1}=m-1$, we have
$\varphi_{N,\rho_{N}(m-1)}(x)=\left(\frac{a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a\big)}\right)^{N-1}\\ \cdot\frac{W_{N-1}(x)\varphi_{N-1,\rho_{N-1}(m-1)}(x)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(m-1-n^{\prime})\big)},$ (3.52)
where we have used the fact that both $W_{N-1}(x)$ and $\varphi_{N-1,\rho_{N-1}(m-1)}$ are invariant under translations $x\to x+(t,\ldots,t)$ for any $t\in{\mathbb{C}}$. Using (3.1) and (3.48) for $N\to N-1$, we easily arrive at the right-hand side of (3.48). ∎
To establish quasi-invariance, we make use of the fact that $\varphi_{N}$ is $S_{N}$-invariant in the sense of the following lemma.
Lemma 3.8.
We have
$\varphi_{N}(\sigma(x),\sigma(y))=(-)^{|\sigma|}\varphi_{N}(x,y),\ \ \ \sigma\in S_{N}.$ (3.53)
Proof.
From (3.36) and the induction assumption, we see immediately that the statement holds true whenever $\sigma\in S_{N-1}$. Hence, we need only to consider the special case $\sigma=\sigma_{N-1}$. By a direct computation, similar to the one in the proof of Proposition 3.7, we obtain
$\begin{split}\varphi_{N,\sigma_{N-1}(\rho_{N}(m-1))}(x)&=-\left(\frac{a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{1}{\prod_{k=1}^{N}\prod_{j=1}^{\min(k,N-1)-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in\big)}\\ &\quad\cdot\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{N-1}-x_{N}+in\big)}\\ &=-\varphi_{N,\rho_{N}(m-1)}(\sigma_{N-1}(x)).\end{split}$ (3.54)
(Note that only the term in (3.49) corresponding to $\nu^{\prime}=\rho_{N-1}$, $n_{1}=\cdots=n_{N-2}=m-1$ and $n_{N-1}=0$ contributes.) It follows that $\varphi_{N}(x,y)$ and $\varphi_{N}(\sigma_{N-1}(x),\sigma_{N-1}(y))$ are both (finite) series of the form
$e^{\frac{2\pi i}{a}(x-i\rho_{N}(m-1),y)}f_{N}(x,y),\ \ \ f_{N}(x,y)=\Gamma_{0}(x)+\sum_{\nu>0}\Gamma_{\nu}(x)e^{-\frac{2\pi}{a}(\nu,y)},$ (3.55)
where $\Gamma_{0}=\varphi_{N,\rho_{N}(m-1)}$ and $\nu>0$ means that $\nu=\sum_{1\leq j<k\leq N}l_{jk}(e_{j}-e_{k})$ with $l_{jk}\geq 0$ and not all equal to zero. From Proposition 3.7, we get that the functions $\varphi_{N}(\sigma(x),y)$, $\sigma\in S_{N}$, form a linearly independent set; and, since the pre-factor $\prod_{j<k}2\sinh(\pi(y_{j}-y_{k}))$ in (3.38) is (anti)periodic under shifts $y_{j}\to y_{j}-i$ and the Macdonald operators $D_{N}^{r}$ are $S_{N}$-invariant, it follows that $\varphi_{N}(x,y)$ and $\varphi_{N}(\sigma_{N-1}(x),\sigma_{N-1}(y))$ satisfy a difference equation
$D_{N}(y)\varphi(x,y)=\varphi(x,y)\cdot\text{const.}\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}},\ \ \ D_{N}(y)\equiv D^{1}_{N}(q^{i2y};q^{2},q^{2(1-m)}).$ (3.56)
For the corresponding series $f_{N}$, this translates into a difference equation $L_{N}(y)f(x,y)=f(x,y)\cdot\text{const.}\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}}$ given by the difference operator
$\begin{split}L_{N}(y)&\equiv e^{-\frac{2\pi i}{a}(x-i\rho_{N}(m-1),y)}D_{N}(y)e^{\frac{2\pi i}{a}(x-i\rho_{N}(m-1),y)}\\ &=\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}}\prod_{k<j}\frac{1-e^{\frac{2\pi}{a}(y_{j}-y_{k}+i(m-1))}}{1-e^{\frac{2\pi}{a}(y_{j}-y_{k})}}\cdot\prod_{k>j}\frac{1-e^{\frac{2\pi}{a}(y_{k}-y_{j}-i(m-1))}}{1-e^{\frac{2\pi}{a}(y_{k}-y_{j})}}\cdot e^{-i\partial_{y_{j}}}.\end{split}$ (3.57)
By expanding the coefficients in power series in $e^{\frac{2\pi}{a}(y_{k}-y_{j})}$, $1\leq j<k\leq N$, this, together with the leading coefficient $\Gamma_{0}$, is readily seen to uniquely determine $f_{N}$. This essentially amounts to the uniqueness result in Theorem 2.3 in [NS12], wherein a detailed proof can be found. ∎
Proposition 3.9.
For all $1\leq j<k\leq N$ and $s=1,\ldots,m-1$, we have the vanishing property
$\varphi_{N}\left(x;y+\frac{is}{2}(e_{j}-e_{k})\right)=\varphi_{N}\left(x;y-\frac{is}{2}(e_{j}-e_{k})\right),\ \ \ e^{\frac{2\pi}{a}(y_{j}-y_{k})}=1.$ (3.58)
Proof.
Thanks to Lemma 3.8, it suffices to establish the vanishing property when $j=1$ and $k=2$; and, in this case, it follows immediately from (3.36) and our induction assumption. ∎
Comparing Props. 3.7 and 3.9 with the properties (2.16) and (2.18)–(2.19), which uniquely characterises the self-dual BA-function $\psi_{N}$, we find the following result.
Proposition 3.10.
Given $m\in\mathbb{Z}_{+}$ and a>m1a>m-1, we have
$\begin{split}\varphi_{N}(a,m;x,y)&=\left(\frac{a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{\psi_{N}(\exp(-i\pi/a),m-1;ix,iy)}{\prod_{1\leq j<k\leq N}\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}.\end{split}$ (3.59)
Finally, by combining Props. 3.6 and 3.10, we arrive at the statement in Thm. 3.1 for arbitrary NN; and, by using the above expression to express $\varphi_{N}$ and $\varphi_{N-1}$ in (3.36) in terms of $\psi_{N}$ and $\psi_{N-1}$, respectively, we easily verify the claim in Prop. 3.2.
Acknowledgements
I would like to thank N. Belousov for helpful comments. Financial support from the Swedish Research Council (Project-id 2024-05649_VR) is gratefully acknowledged.